Prev Next

SW&IT Meeting Minutes 2020-11-09

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20201102
LR20201102 - Bec, Matthieu 40.0
LR20201102 - Holmberg, Carl 32.0
LR20201102 - Edmeades, David 40.0
LR20201102 - Sola, Igor 0.0
LR20201102 - Grenz, Paul 20.0
LR20201102 - Hooper, Stephen 80.0
LR20201102 - Bryant, Trey 0.0
LR20201102 - Zhang, Yang 40.0
LR20201102 - Zhang, Xianyu 0.0

News from management

  • Veterans Day holiday

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT
    • DMS GS oMag from tcs logs
    • cobbler pxe/kickstart
  • This week
    • ATT
    • cobbler pxe/kickstart recap
    • DMS soul-telemetry

DE

  • Sorted out backup VMs
    • Will set autologin
  • Archive machine stuff
    • Reminder to archive team re: archive.mountain
    • Took usage logs of archive; will plot against instrument usage
  • L4AO site stuff
    • Got security reminder; we don't use anything complex
    • Decided not to bother with SSL cert
  • Account creations for VPN and IssueTrak
Next:
  • Syno support tasks
  • Office trip, planning Tuesday evening
    • UPS batteries
    • UPS balancing
    • General cleanup
    • Verified negative RDP connection to Gene's PC; will check its status
  • Ping LD about drive swap
  • Cisco training

CH

  • Last week:
    • AGW: Investigated 'pmac' client firmware upload to PMAC. May not be possible.
    • DIMM: Added package reengineering issues to github.
    • statserv:
      • Discovered the INDI service was down, leaving the AHS status page w/o updates.
      • Patched the 'runindi' service script for more consistent process testing, lower period CPU loading.
    • TCS:
      • With MB, SH, YZ, discovered and disabled unneeded TCS:DDS:ICE logging flood on tcs2.
      • Disabled dev lbto-ahs-ioc.service, which was causing significant spurious logging on tcs2.
    • Safety and harrasment training.
  • This week:
    • AGW:
      • Capture new PMAC 'firmware' from JP in github
      • Contact Delta Tau to revisit 'pmac' upload capability.
      • Test encoder in-situ to check the value is correct (rotating the encoder by N degrees)
    • AHS: Fix byte stream to string conversion error that has popped back up. Possibly caused by change in LBTI INDI parameters.
    • statserv: investigate and test which INDI drivers, if any are explicitly required, to restore web page updates.

YZ (may not join the meeting due to TMS test)

  • Previous:
    • TMS work (discusson/meeting and data analysis, new tests planned).
    • AO-Cravat/OSCO work (some additional software work plus meetings for new plans).
    • Some minor TCS/LBC work.
    • Requested UA harassment training.
  • This week:
    • New TMS tests and continuing software work.
    • AO-Cravat/OSCO software work.
    • modsAlign refactor.
    • Some TCS/LBC work (AOS patch, DDS change for future release, LBC repository reorganization, etc.).
    • Hopefully finish the UA harassment training.

PSG

  • Last week:
    • Working on reading and understanding the documentation for the SIEMANS D425 controller.
  • This week:
    • Working on outlining and starting coding a command-line based program to command this controller.

IS

Last week:
  • Harassment training
  • LUCI
    • Geirs script
    • Started looking into RPMs generation
    • IT 8231 Geirs timeout
  • PIT
    • pit admin tool: good progress on proposal listing table, and backend functionality. (On going). i.e. generating and uploading new pdfs, sending email, populating dbs etc.

This week:
  • Harassment training
  • PIT
    • pit admin tool
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.

SH

  • IP addresses, DNS JP
  • Laura pop-ups
  • Wired MAC submitted, accepted.
  • Satellite image -- need to fix this all.
  • Jim and PC
  • LUCI, robs
  • cobbler VM (not needed)
  • Synology and Google LDAP (no good)
  • Power meeting with Steward.

next:

  • Order laptop, monitors
  • termserv1
  • Synology AD replacement
  • Wiki sync

XZ

  • Last week
    • AOS audio for AO loop open
    • ATT AdSec
    • ATT SW
    • “all processes disconnected from Msgd and idlcltr NOT running” issue (IT 8217)
    • AO operation support
  • This week
    • ATT AdSec
    • ATT SW
    • AOS patch on the mountain
    • SX AdSec retracted test and cause study
Topic revision: r19 - 09 Nov 2020, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback